If you note the following yield curve in The Wall Street Journal, what is the one-year forward rate for the period beginning one year from today, 2f1 according to the unbiased expectations theory? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16)) Maturity Yield One day 2.56 % One year 2.68 Two years 2.82 Three years 2.93
Answer #1

If unbiased expectations theory holds, then :

(1 + 1 year rate) * (1 + 2f1) = (1 + 2 year rate)2

(1 + 2.68%) * (1 + 2f1) = (1 + 2.82%)2

2f1 = (1.02822 / 1.0268) - 1

2f1 = 2.96%

The one-year forward rate for the period beginning one year from today is 2.96%